​(Capital structure​ analysis)  The liabilities and​ owners' equity for Campbell Industries is found​ here:  

.

Accounts payable

$ 459 comma 000$459,000

Notes payable

$250,000

Current liabilitie

$709,000

​Long-term debt

$1,217,000

Common equity

$4,841,000

Total liabilities and equity

$6,767,000

a.  What percentage of the​ firm's assets does the firm finance using debt​ (liabilities)?

b.  If Campbell were to purchase a new warehouse for

$ 1.3$1.3

million and finance it entirely with​ long-term debt, what would be the​ firm's new debt​ ratio?

a.  What percentage of the​ firm's assets does the firm finance using debt​ (liabilities)?

The fraction of the​ firm's assets that the firm finances using debt is

nothing​%.

​(Round to one decimal​ place.)
Answer #1

(a) Total Liabilities= Current Liabilities + Long term Debt = $709,000 +$1,217,000= $1,926,000

Total assets equal to total of liabilities and equity= $6,767,000

Fraction of assets financed by debt= (Liabilities/Total assets)*100 = ($1,926,000/$6,767,000)*100= 28.461652%

(b) Cost of new warehouse= $1,300,000 (financed with long term debt)

Total assets with the new warehouse= $6,767,000 + $1,300,000 = $8,067,000

Total liabilities (debt) = $1926000 + $1,300,000 = $3,226,000

New Debt Ratio= (Total Liabilities/ Total Assets)*100 = ($3,226,000/$8,067,000)*100 = 39.990083%